OpenStudy (anonymous):

What is the general form of a polynomial p(x) such that that\[ p(x^2)=p(x)^2\]

OpenStudy (anonymous):

x^n

OpenStudy (anonymous):

Darn, I went to take a drink of OJ and Avinash wrote it first.

OpenStudy (anonymous):

\[f(x)=x\]

OpenStudy (anonymous):

lol ok i l make up for that

OpenStudy (anonymous):

You're too nice, Avinash! :)

OpenStudy (anonymous):

thanks i guess :)

OpenStudy (anonymous):

or rather \[\left( f(x)\right) ^n=x^n\]

OpenStudy (anonymous):

x^n

OpenStudy (anonymous):

@eliassaab how do we can prove that only root of polynomial is 0?

OpenStudy (anonymous):

Apart from 0 complex numbers with |a| = 1 are only possible roots but i cant cancel them. it seems easier than the other problems but i dony know...

OpenStudy (anonymous):

i think its trivial because there r infinite complex numbers with |a| = 1 and if one of them is root then the others must be too and this is not true because a polynomial can only have a finite number of roots

OpenStudy (anonymous):

@eliassaab thank u sir for the nice problems u put up